# EliGen: Entity-Level Controlled Image Generation with Regional Attention

### v1

@@ 3.4 Implementation Details
Our model does not introduce any additional parameters. To achieve faster convergence and better alignment with the community, we utilize the LoRA (Low-Rank Adaptation) Hu et al. (2021) method for fine-tuning. The LoRA weights are applied to the linear layers of each block in DiT, including the projection layers before and after Attention, as well as the linear layers within Adaptive Layer Norm. The LoRA Rank is set to 64. The model is trained with a batch size of 64 for 20K steps using the AdamW optimizer and a learning rate of 0.0001. To minimize memory usage, we employ the DeepSpeed Rasley et al. (2020) training strategy with a bfloat16 precision. The model is trained with the standard generation loss of rectified flow Esser et al. (2024). During inference, we utilize the Euler scheduler Karras et al. (2022) with 50 sampling steps and set the classifier-free guidance (CFG) scale to 3. Additionally, the embedded guidance within Flux is configured to 3.5.
4 Application
@@
Figure 6: Qualitative results conditioned on arbitrary-shape entities.
Figure 7: Qualitative comparison on image inpainting. Global prompt for all methods are generated by MLLM. For multi-entity inpainting with ControlNet, the result is achieved through multiple sequential forward passes.
@@ 4.1 Entity-Level Controlled Generation
In this scenario, entities are localized using bounding boxes. Our model converts bounding boxes to masks for input, while others utilizer bounding box coordinates. The qualitative results are shown in Figure 5, where our model demonstrates superior performance in image quality, coherence and entity accuracy. In the first two rows, simple tasks with distinct entities are handled by most methods except Multi-Diffusion. The last two rows highlight complex layouts where entities are intricately coupled. While other models struggle to maintain accuracy and coherence, even MIGC and Multi-Diffusion falter in the final example, our model consistently delivers high-quality, coherent images.
@@
Prior to this work, Instance Diffusion was the sole approach capable of handling arbitrary-shaped masks as input, prompting us to conduct a comparative evaluation, as illustrated in Figure 6. Instance Diffusion exhibited poor image quality when relying solely on the mask, so we augmented it with the mask’s minimal bounding box as additional input. The results clearly show that our model outperforms Instance Diffusion in image quality, positional accuracy, and attribute fidelity. Instance Diffusion struggles with position control for masks that significantly deviate from rectangular shapes, such as “Chinese dragon” and “Great Wall”. In contrast, our model, without hardcoded positions, consistently delivers excellent results.
@@ 4.2 Image Inpainting
Image inpainting targets rendering desired content at specified image locations. While encoding the input image as the initial latent for the EliGen model is intuitive, it often disrupts non-inpainting regions. Thus, we propose the inpainting fusion pipeline to preserve these areas while enabling precise, entity-level modifications over inpaining regions.
Given image $I$, we first encode it as embedding $z_{init}$ with a VAE encoder. Then, for each denoising step $t$, we compute the desired noise of inpainting foreground $Nf_{t}$ and background $Nb_{t}$ as follows.
$Nf_{t}$ $=\mathrm{DiT}(z_{t};p;{p_{1},\dots,p_{n}};{m_{1},\dots,m_{n}})$ (14)
$Nb_{t}$ $=(z_{t}-z_{init})/\sigma_{t}$ (15)
where $m_{i}$ is the inpainting mask of the $i_{th}$ entity, $z_{t}$ denotes the latent embedding needed denoising and $\sigma_{t}$ is a hyperparameter controlling the noise strength. Then, the noise is calculate by fusion the foreground of $Nf_{t}$ and the background of $Nb_{t}$.
$Npos_{t}=Nf_{t}\cdot\cup m_{i}+Nb_{t}\cdot(1-\cup m_{i})\$ (16)
To enabling classifier-free guidance, we further compute the negative noise $Nneg_{t}$ and the final noise $N_{t}$ as follows with a cfg scale $cfg$:
$Nneg_{t}=\mathrm{DiT}(p_{neg})$ (17)
$N_{t}=Nneg_{t}+cfg\cdot(Npos_{t}-Nneg_{t})$ (18)
Finally, using $z_{t}$ and $N_{t}$, we can obtain the denoised embedding $z_{t-1}$ with the diffusion scheduler:
$z_{t-1}=z_{t}+N_{t}\cdot(\sigma_{t-1}-\sigma_{t})$ (19)
We compare our method with Flux-ControlNet Creative (2024) in Figure 7. As is shown in the figure, inpainting fusion ensures background invariance (row 2), and training regional attention further enables EliGen to render desired objects in the inpainting region (row 3). Compared to Flux-ControlNet (row 1), our model excels at: (1) preserving non-inpainting areas without quality loss, (2) enabling precise control over entity position and shape for harmonious blending, and (3) supporting multi-entity inpainting in a single forward pass.
Figure 8: Styled entity control with IP-Adapter and our EliGen.
Figure 9: Dialogue based image design and editing with MLLM and our EliGen.
@@ 4.3 Potential Creative Extension
EliGen holds boundless creative potential, and here we demonstrate its capabilities when integrated with IP-Adapter and MLLM.
Styled Entity Control with IP-Adapter
IP-Adapter Team (2024) enables the generation of target images based on the style of a reference image. When combined with EliGen, it allows us to accomplish the styled entity control task. Figure 8 illustrates some examples along with their generated variants. By inputting multiple entity details alongside a reference image, we achieve precise control over both the characters and the environmental style of the target image.
Interactive Image Generation and Editing with MLLM
The MLLM is capable of interpreting user-provided images and textual content to engage in dialogue. As a result, we can achieve interactive image design and editing by leveraging the integration of MLLM and EliGen. Specifically, grounding-enabled MLLMs such as Qwen2-VL can be utilized to determine entity locations based on textual descriptions, enabling the generation of images solely through text using EliGen. Furthermore, the MLLM can be employed to comprehend images and design entity information in accordance with specific image editing requirements, thereby facilitating precise image editing in conjunction with EliGen. The qualitative results are shown in Figure 9.
5 Conclusion and Future works
@@
To achieve image generation with entity controls, we propose regional attention into the DiT architecture to effectively integrate entity prompts and their locations. Furthermore, we construct a dataset with fine-grained entity information and finetune the regional attention mechanism, developing the EliGen model. EliGen demonstrates high generalization in entity-level controlled image generation, delivering superior image quality compared to existing models. We further propose an inpainting fusion pipeline, extending EliGen’s functionality to image inpainting tasks. Moreover, by integrating EliGen with community models such as IP-Adapter and MLLM, we demonstrate its potential for creative applications, broadening its scope and versatility.
EliGen opens promising avenues for future research. We will further explore the integration of additional control modalities, such as depth and canny maps, to improve synthesis precision. Advanced generation paradigms, including grouped diffusion via in-context LoRA and style-specific LoRA models, will be investigated to enhance output diversity. Its unique ability to utilize arbitrary-shaped masks for high-quality generation positions it as a powerful tool for large-scale, customized data synthesis, with potential applications in synthetic datasets, virtual reality, and beyond. Collaborative, community-driven innovation will further unlock its potential for both creative and practical applications.

### v2

@@ 3.4 Implementation Details
Our model does not introduce any additional parameters. To achieve faster convergence and better alignment with the community, we utilize the LoRA (Low-Rank Adaptation) Hu et al. (2021) method for fine-tuning. The LoRA weights are applied to the linear layers of each block in DiT, including the projection layers before and after attention, as well as the linear layers within adaptive LayerNorm. The LoRA Rank is set to 64. The model is trained with a batch size of 64 for 20K steps using the AdamW optimizer and a learning rate of 0.0001. To minimize memory usage, we employ the DeepSpeed Rasley et al. (2020) training strategy with a bfloat16 precision. The model is trained with the standard generation loss of rectified flow Esser et al. (2024). During inference, we utilize the Euler scheduler Karras et al. (2022) with 50 sampling steps and set the classifier-free guidance Ho and Salimans (2022) scale to 3. Additionally, the embedded guidance within FLUX is configured to 3.5.
4 Qualitative Experiments
@@
Figure 5: Qualitative results with arbitrary-shaped entities.
@@ 4.1 Entity-Level Controlled Generation
In this scenario, entities are localized using bounding boxes. Our model converts bounding boxes to masks for input, while others utilize bounding box coordinates. The results are shown in Figure 4, where our model demonstrates superior performance in image quality, and entity accuracy. In the first two rows, simple tasks with distinct entities are handled by most methods except Multi-Diffusion. The last two rows highlight complex layouts where entities are intricately coupled. While other models struggle to maintain accuracy and coherence, our model consistently delivers high-quality, coherent images.
@@
We further conduct experiments to verify the generalization ability of our method for arbitrary-shaped masks. Prior to this work, Instance Diffusion was the sole approach capable of handling arbitrary-shaped masks as input. The results, illustrated in Figure 5, clearly show that EliGen outperforms Instance Diffusion in image quality, positional accuracy, and attribute fidelity. Instance Diffusion struggles with position control for masks that significantly deviate from rectangular shapes, such as “Chinese dragon” and “Great Wall”. In contrast, EliGen consistently delivers excellent results.
@@ 4.2 Image Inpainting
Image inpainting targets rendering desired content at specified locations. While encoding the input image as the initial latent for the EliGen model is intuitive, it often disrupts non-inpainting regions. Thus, we propose the inpainting fusion pipeline to preserve these areas while enabling precise entity-level modifications over inpaining regions. The core mechanism involves integrating entity latents from inpainting regions with background latents from non-inpainting areas. Due to space constraints, the algorithmic details are provided in the Supplementary Material.
We compare our method with FLUX-ControlNet Creative (2024) in Figure 6. As is shown in the result, inpainting fusion ensures background invariance, and training regional attention further enables EliGen to render desired objects in the inpainting region. Compared to FLUX-ControlNet, our model excels at: (1) preserving non-inpainting areas without quality loss, (2) enabling precise control over entity position and shape for harmonious blending, and (3) supporting multi-entity inpainting in a single forward pass.
Table 1: Evaluation on COCO benchmark from multiple dimensions: Entity Success Rate, Spatial Accuracy and Image Quality. $\uparrow$ indicates that a higher value is better. Bold and underline denote the best and second best methods, respectively.
Model Entity Success Rate (%)
Spatial
Accuracy (%)
Image Quality
Qwen2VL $\uparrow$ InternVL2 $\uparrow$ CogVLM2 $\uparrow$ mIoU $\uparrow$
Entity
CLIP $\uparrow$
Global
CLIP $\uparrow$
Aesthetic
Score $\uparrow$
Multi-Diffusion 51.17 50.38 83.15 31.74 20.07 22.58 5.04
GLIGEN 71.83 68.47 88.51 59.48 20.15 24.94 5.07
MIGC 77.01 74.48 89.76 64.75 20.48 23.57 5.06
Instance Diffusion 82.43 79.49 91.33 76.17 21.06 24.17 4.89
EliGen (Ours) 88.41 83.79 96.25 73.93 21.47 27.39 5.60
Table 2: Human preference evaluation on COCO benchmark. The numerical values represent the win rate in pairwise comparisons. All scores significantly exceeding 50% demonstrate that EliGen aligns more closely with human preferences compared to other methods.
Comparison MPS (%) PickScore (%)
EliGen vs. Instance Diffusion 71.19 65.70
EliGen vs. Gligen 60.71 60.55
EliGen vs. MIGC 74.44 69.48
EliGen vs. Multi-Diffusion 69.60 72.32
@@ 4.3 Potential Creative Extension
EliGen holds boundless creative potential, and we demonstrate its capabilities when integrated with IP-Adapter Team (2024), In-Context LoRA Huang et al. (2024a) and MLLM Wang et al. (2024a). Due to space constraints, we provide detailed methods and results in the Supplementary Material, and we highly recommend reviewing them.
Figure 6: Qualitative comparison on image inpainting. Global prompt for all methods are generated by MLLM.
5 Quantitative Experiments
5.1 Experimental Setup
Following previous methods, we perform quantitative evaluations on the filtered COCO Lin et al. (2014); Zheng et al. (2023) 2017 validation set, comprising 2830 prompt-image pairs. Previous GLIGEN Li et al. (2023) utilized the mean average precision for evaluation, a metric that computes the area under the precision-recall curve. However, we observe its critical limitation: it treats background entities in generated images as negative samples, resulting in substantial inaccuracies. As a result, we decided not to use this metric. Instead, to thoroughly evaluate the models’ capabilities on the COCO benchmark, we assess their performance across the following dimensions.
Entity Success Rate
To evaluate the model’s ability to generate entities accurately, we assess the quality of the target regions, entity details, and image-text alignment with MLLMs. The success rate is then calculated by aggregating results across all entities in the benchmark.
Spatial Accuracy
We evaluate the spatial accuracy of generated entities by calculating the mean Intersection-over-Union (mIoU) between the ground truth bounding boxes and the predicted bounding boxes. The predicted bounding boxes are obtained using YOLOv11x Khanam and Hussain (2024).
Image Quality
We evaluate semantic accuracy using the CLIP score for both entities and the global image. Additionally, Aesthetic Score is adopted to measure the visual quality.
Human Preference
To assess the alignment with human preferences, we compare EliGen against other models using two metrics: the Multi-dimensional Preference Score (MPS) Zhang et al. (2024) and PickScore Kirstain et al. (2023).
5.2 Comparison on COCO Benchmark
The evaluation results of the models on the COCO benchmark are presented in Tables 1. For Entity Success Rate, we assessed all methods using Qwen2-VLWang et al. (2024a), InternVL2Chen et al. (2024), and CogVLM2 Hong et al. (2024). The results demonstrate that EliGen consistently outperformed other models, highlighting its robustness and high quality. In terms of mIoU, our model performs comparably to Instance-Diffusion and substantially better than other methods. The slightly lower mIoU of EliGen compared to Instance-Diffusion is primarily due to its soft-coding scheme for positional information. This scheme allows EliGen to automatically expand certain regions to ensure image coherence, as further illustrated in the Supplementary Material. In terms of both CLIP Score and Aesthetic Score, EliGen significantly outperforms other models, demonstrating substantial superiority in image quality.
The human preference results in Table 2 measure the probability of humans favoring one image over another in pairwise comparisons. EliGen consistently achieves a preference probability well above 50% against all other models, demonstrating its stronger alignment with human preferences and higher image quality of generated images.
Table 3: Results of user study. Each user was asked to rate an image across the four dimensions on a scale from 1 to 5.
Model
Spatial
Accuracy
Entity
Details
Entity
Coherence
Image
Aesthetic
Multi-Diffusion 2.90 2.49 2.53 2.38
GLIGEN 4.23 3.24 3.53 3.28
MIGC 4.15 2.82 3.09 2.79
Instance Diffusion 4.45 3.37 3.51 3.31
EliGen (Ours) 4.62 4.44 4.31 4.30
5.3 User Study
To complement the automated metrics, we conducted a user study for all models. We randomly select 250 samples with $1\sim 3$ entity conditions from the COCO benchmark and invite users to evaluate these examples across multiple dimensions, including Spatial Accuracy, Entity Details, Entity Coherence, and Image Aesthetic. The Entity Coherence metric assesses how well the entities integrate with other entities or the environment, ensuring that no isolated entities appear in the generated images.
We collected a total of 5,500 ratings, with the evaluation results presented in Table 3. EliGen achieves the highest scores across all four dimensions, particularly excelling in Entity Details, Entity Coherence, and Image Aesthetic. These results are consistent with those in Table 1, demonstrating EliGen’s robustness in controlling entities with higher spatial precision and superior image quality.
6 Conclusion
@@
To achieve image generation with entity-level controls, we propose regional attention into the DiT architecture to effectively integrate entity prompts and their locations. Furthermore, we construct a dataset with fine-grained entity information and finetune the regional attention mechanism, developing the EliGen model. EliGen demonstrates high generalization in entity-level controlled image generation, delivering superior image quality and entity accuracy compared to existing models. We further propose an inpainting fusion pipeline, extending EliGen’s functionality to image inpainting tasks. Moreover, by integrating EliGen with open-source models, we demonstrate its potential for creative applications, broadening its scope and versatility.
Supplementary Material
Appendix A Dataset Details
A.1 Dataset Construction Pipeline
Figure 7: The dataset generation and annotation pipeline. We utilize Flux to generate images and annotate them with Qwen2-VL.
As illustrated in Figure 7, to circumvent the additional training overheads that may arise from data distribution differences, we leverage FLUX.1-dev Labs [2024] for the direct generation of training images. Following this, we deploy Qwen2-VL 72B Wang et al. [2024a] to annotate the images. This process involves crafting a comprehensive global description of the image, identifying entities present within the image, and meticulously annotating each entity with textual descriptions and their respective spatial locations. Ultimately, we create approximately 500K training samples with 1.27M entity annotations.
Figure 8: The dataset distributions of entity numbers and categories.
A.2 Dataset Statistics
We conduct a statistical analysis of the annotated dataset in Figure 8, primarily focusing on the distribution of the number of entities per image and the semantic categories of the entities.
In terms of the number of entities, the vast majority of samples contain 1 to 5 entities, while approximately 2% of the samples have more than 6 entities. This distribution of entities is training-friendly because a stable number of entities ensures that the memory usage and computational resource consumption per training step remain relatively consistent. Additionally, since our designed regional attention mechanism supports a variable number of entities, the trained EliGen model demonstrates strong generalization capabilities even for cases with more than 6 entities, effectively handling multi-entity scenarios.
Regarding entity categories, our dataset demonstrates substantial diversity, with entities broadly classified into 7 categories. Since annotations are open-set (not limited to predefined categories), we use Qwen2.5 3B Yang et al. [2024] to categorize them into the following groups for statistical analysis:
  • Person: human beings, including men, women, children, and fictional characters.
  • Scenery: natural landscapes, such as mountains, forests, rivers, and beaches.
  • Architecture: buildings, structures, and man-made environments.
  • Clothing: apparel and accessories worn by people.
  • Animal: living creatures, including pets and wild animals.
  • Everyday Objects: common items used in daily life, such as furniture, tools, and gadgets.
  • Other: all entities that do not fit into the above categories, including fictional elements.
Appendix B Inpainting Fusion Algorithm
Algorithm 1 Image Inpainting Fusion Algorithm
Input: Input image $I$, inpainting masks $\{m_{1},\dots,m_{n}\}$, global prompt $p$, local prompts $\{p_{1},\dots,p_{n}\}$, negative prompt $p_{neg}$, and hyperparameters $\{\sigma_{t}\}_{t=1}^{T}$, $cfg$; VAE model $\mathrm{VAE}=(E(I),D(z)$} and Diffusion Transformer $\mathrm{DiT}(z,p,m,t)$}
Output: Inpainted image $P$ with entities $\{p_{1},\dots,p_{n}\}$ inpainted on regions $\{m_{1},\dots,m_{n}\}$
   Encode the input image: $z_{init}\leftarrow E(I)$
   Initialize $z_{T}$ with random noise: $z_{T}\sim\text{noise}(z_{init},T)$
   for each denoising step $t=T,\dots,0$ do
      Compute foreground noise: $Nf_{t}\leftarrow\mathrm{DiT}(z_{t};p;\{p_{1},\dots,p_{n}\};\{m_{1},\dots,m_{n}\};t)$
      Compute background noise: $Nb_{t}\leftarrow(z_{t}-z_{init})/\sigma_{t}$
      Fuse foreground and background noise: $Npos_{t}\leftarrow Nf_{t}\cdot\cup m_{i}+Nb_{t}\cdot(1-\cup m_{i})$
      Compute negative noise: $Nneg_{t}\leftarrow\mathrm{DiT}(z_{t};p_{neg};t)$
      Apply classifier-free guidance: $N_{t}\leftarrow Nneg_{t}+cfg\cdot(Npos_{t}-Nneg_{t})$
      Update latent embedding with flow matching pipeline: $z_{t-1}\leftarrow z_{t}+N_{t}\cdot(\sigma_{t-1}-\sigma_{t})$
   end for
   Decode the denoised latents: $P=D(z_{0})$
   Return: Inpainted image $P$
Figure 9: Styled entity control with IP-Adapter and our EliGen. Upon receiving the desired entity information, we can generate a target image that mirrors the style and characters of the reference image, while ensuring the position and attributes of the entities are precisely controlled.
Image inpainting aims to reconstruct or modify specific regions of an input image, such as adding or replacing objects. To leverage the entity-level generation capabilities of EliGen for image inpainting while ensuring that non-inpainting regions remain unchanged and the inpainting regions align with the overall image style, we propose the Inpainting Fusion Algorithm, as detailed in Algorithm 1. The input image $I$ is first encoded into an initial latent embedding $z_{init}$. At each denoising step, the predicted noise from the DiT is fused with the noise derived from the original image. This fusion ensures that the foreground regions are inpainted with the generated entities, while the background regions remain consistent with the original image. Furthermore, we incorporate classifier-free guidance to refine the overall image quality during the fusion process. After $T$ denoising steps, the denoised latent embedding $z_{0}$ is decoded to produce the final inpainted image $P$. This approach achieves precise entity-level modifications while maintaining the coherence and stylistic consistency of the entire image.
Figure 10: Entity transfer with In-Context LoRA and our EliGen. Given the identity defined by the source image, EliGen can seamlessly transfer it to any desired location and scale within the target image, thereby achieving the ideal visual design.
Figure 11: Dialogue-based image design and editing with MLLM and our EliGen. Users can utilize MLLM to design the layout of the image they wish to generate or input the image they want to edit into MLLM. After MLLM generates the caption and entity information of the image, EliGen can produce the target image.
Appendix C Potential Creative Extension
EliGen holds boundless creative potential, and here we demonstrate its capabilities when integrated with IP-Adapter Team [2024], In-Context LoRA Huang et al. [2024a] and MLLM Wang et al. [2024a].
Figure 12: Training qualitative comparison between our constructed dataset and Laion dataset. The image pair is generated from the same noise, and the global prompt is “A person standing by the river.” The comparative results show that the images generated by the model trained on our proposed dataset significantly outperform those from the Laion dataset in quality.
C.1 Styled Entity Control with IP-Adapter
IP-Adapter Team [2024] enables the generation of target images based on the style of a reference image. When combined with EliGen, it allows us to accomplish the styled entity control task. Figure 9 illustrates some examples along with their generated variants. By inputting multiple entity details alongside a reference image, we achieve precise control over both the characters and the environmental style of the target image.
C.2 Entity Transfer with In-Context LoRA
In-Context LoRA is capable of performing few-shot image composition tasks conditioned on either global prompts or images, and can be applied in various scenarios such as portrait photography, product design, font design, and visual identity design. Our EliGen model further enhances these scenarios by incorporating entity-level control conditions, enabling more personalized image generation. Taking visual identity design as an example, we have successfully accomplished the entity transfer task by introducing EliGen, as illustrated in Figure 10. Given the entity defined by the source image, we can transfer it to any desired location in the target image through EliGen’s positional control capability, thereby achieving the intended visual identity.
C.3 Dialogue-based Image Design and Editing with MLLM
The MLLM is capable of interpreting user-provided images and textual content to engage in dialogue. As a result, we can achieve dialogue-based image design and editing by leveraging the integration of MLLM and EliGen. Specifically, grounding-enabled MLLMs such as Qwen2-VL can be utilized to determine entity locations based on textual descriptions, enabling the generation of images solely through text using EliGen. Furthermore, the MLLM can be employed to comprehend images and design entity information in accordance with specific image editing requirements, thereby facilitating precise image editing in conjunction with EliGen. The qualitative results are shown in Figure 11.
Appendix D Ablation Study
D.1 Training Qualitative Comparison for Dataset Image Source
The primary objective of training the EliGen model is to activate its capability for layout control, which necessitates the construction of a dataset annotated with entities. Moreover, the image distribution within the dataset should closely align with the output distribution of the model itself. If the distribution gap is substantial, a significant portion of the training loss would be attributed to this discrepancy. This would markedly impede the model’s convergence and degrade the quality of the generated images. Consequently, we opt to utilize the Flux-generated images for dataset creation, rather than directly employing images from the open-source Laion dataset LAION [2022].
We conduct training experiments using both our custom-created dataset and the Laion dataset, where the annotations for the Laion dataset a re s derived from the same pipeline described in Section A.1. As illustrated in Figure 12, under identical prompts and initial noise conditions, the model trained on our dataset demonstrates significantly superior image quality compared to the model trained on the Laion dataset. Compared to Flux-generated images, the latter exhibit significant visual clutter, inferior detail quality, and a notable deviation from the expected distribution.
Figure 13: Attention maps of the local prompt “person” with latent. The visualization results are derived from the regional attention (RA) layer at the last double-stream transformer block, showing the first ten denoising timesteps. As denoising progresses, the attention maps gradually reveal the shape of the entity, and the activation regions of the trained RA are primarily concentrated within the target control area.
Figure 14: Generalization tests across different random seeds. Variations in the random seed induce changes in the initial noise input to the model. The two examples above demonstrate that EliGen possesses the capability to rectify images to the target layout for arbitrary noise inputs, while simultaneously preserving sufficient image diversity.
D.2 Analysis on Attention Map
Regional Attention (RA) primarily guides the generation of entities in specified regions through attention maps. Therefore, in Figure 13, we visualize the attention maps to demonstrate that RA indeed plays a role in guiding entity generation. Note that here we visualize the attention maps of the entire latent image. In practice, during regional attention, the uncontrolled regions of the attention map are deactivated.
From the visualization results, it can be observed that the trained RA, compared to the untrained one, confines attention activation regions to the controlled area more rapidly. Meanwhile, as denoising progresses, the entity’s shape gradually emerges within the controlled region, while activation in uncontrolled regions diminishes, thereby completing the controlled entity generation.
Appendix E Additional Qualitative Results
EliGen is a robust entity-level control model, exhibiting strong generalization capabilities across diverse positional inputs, semantic contexts, and random noise. In this section, we present additional qualitative results to further validate its robustness.
E.1 Various Random Seeds
Variations in random seeds lead to changes in the initial noise input to the model. For conventional text-to-image models, the layout of the image is typically determined by the initial noise. In contrast, EliGen possesses the capability to control the layout, ensuring that each controlled entity is placed in the target position even when the initial noise varies. Test examples are shown in Figure 14, where each image is generated with different random seeds, yet all adhere to the target layout, demonstrating the generalization ability of EliGen. Notably, the generated samples also exhibit diversity, as variations in random seeds result in changes in the details of each entity.
E.2 Continuously Varying Entity Positions
In Figure 15, we evaluated the control efficacy of EliGen using identical input noise and the same entity. The position of the “person” entity shifts progressively from left to right, with EliGen maintaining consistent control throughout the process. To preserve the identity of the person during generation, we integrate the IP-Adapter Team [2024] method. This demonstrates EliGen’s potential to generate coherent image sequences, such as those involving continuous movement of entities through positional control or complex actions orchestrated via coordinated control of multiple entities.
E.3 Unreasonable Positional Inputs
A robust model must demonstrate generalization capabilities even for suboptimal inputs. For EliGen, these suboptimal inputs primarily pertain to entity positions, which we categorize into two types: incorrect spatial relationships between entities and incorrect entity shapes.
Incorrect Spatial Relationships Between Entities
incorrect spatial relationships between entities refer to scenarios where the relative positions of entities are inaccurate, such as being too far apart or too close together. We evaluated two cases of this inputs, as depicted in Figure 16. In both examples, the correct positioning of the “person” entity necessitates an overlap with another entity to accurately convey the action being performed. For example, in the prompt “A person is playing tennis,” the “person” should ideally be holding a “tennis racket”. However, the input provided to EliGen is incorrect, which is featured two disjointed positions. In such scenarios, EliGen prioritizes the coherence of entities within the image, adaptively fine-tuning their actual positions, such as adjusting the position of the person’s hand. Though this adaption may result in the model underperforming on quantitative metrics related to spatial accuracy, it substantially improves the overall image quality.
Incorrect Entity Shapes
Incorrect entity shapes refer to cases where the input region of an entity does not match its described shape, such as a “standing person” typically being represented by a vertical rectangle shape. Through experimentation, we find that EliGen possesses the generalization to adaptively adjust entity posture based on arbitrary input shape. For example, in Figure 17, EliGen adjusts the gesture of the generated “person” entity according to various positional inputs while maintaining coherence with the environment.
E.4 Pure Text-to-Image Generation Capability
The images generated by EliGen exhibit high quality. When provided with any number of entity inputs, it can achieve precise positional control and detailed generation for each entity. In the absence of any entity inputs, EliGen also functions as an exceptional text-to-image model. As shown in Figure 18, the images generated solely from global prompts demonstrate excellence in lighting, detail, and overall quality.
Figure 15: Continuous variation tests on positional inputs. The samples are generated using the same prompt and initial noise. EliGen can achieve continuous movement of entities, which highlights the potential of EliGen in generating coherent image sequences. To preserve the identity of the “person” during generation, we integrate EliGen with the IP-Adapter method.
Figure 16: Generalization tests with incorrect spatial relationships between entities. In the provided examples, the input positional relationships between entities are inaccurate. For instance, the entities ”person” and ”tennis racket” should intersect rather than being separated to reflect the relationship of a person holding a tennis racket. Under such cases, EliGen prioritizes ensuring the coherence of entities in the image, thereby adaptively fine-tuning the actual positions of the entities, such as adjusting the position of the person’s hand.
Figure 17: Generalization with incorrect entity shapes. EliGen demonstrates the ability to adaptively adjust entity morphology based on input shapes. In the examples, the control shape for the “person” is not a standard upright stance. EliGen adjusts the person’s posture according to the input mask to generate coherent entities.
Figure 18: Pure text-to-image generation capability of EliGen. Despite fine-tuning on an entity-level control task, EliGen retains strong text-to-image generation performance without entity conditions, demonstrating that training does not degrade the model’s image generation quality.